Upon graduating with an accounting degree, you open your own accounting firm of which you are the sole employee. To start the firm you passed on a job offer with a large accounting firm that offered you a salary of RM60,000 annually. Last year you earned a total revenue of RM100,000. Rent and supplies last year were RM50,000. 

a) calculate your annual economic costs,

b) calculate your annual economic profit, and

c) calculate your annual accounting profit.


1
Expert's answer
2021-01-15T09:47:05-0500

a) your annual economic costs are:

TEC = 50,000 + 60,000 = RM110,000.

b) your annual economic profit is:

EP = TR - TEC = 100,000 - 110,000 = -10,000.

c) your annual accounting profit is:

AP = TR - TC = 100,000 - 50,000 = RM50,000.